given A=120m @ 20 degrees W of N and B= 20m @ 10 degrees S of W use the analythic method to find R=A+B
1
Expert's answer
2012-10-08T10:03:37-0400
R=A+B
That means, that if we set projections Rx=X, Ry=Y, by the analitical method (when x from W to E, y from S to N)

X=Ax+Bx

Y=Ay+By

Ax=-A*sin20=-41

Ay=A*cos20=113

By=-B*sin10=-3.5

Bx=-B*cos10=-19.7

So,

Y=113-3.5=109.5

X=-60.7




R=125.2